OpenStudy (anonymous):

2 2/3 divided by 1 1/2

OpenStudy (anonymous):

1.777777777777777333

OpenStudy (anonymous):

9/3 times 3/2 when dividing fractions, you keep the first and flip the second and multiply. called change, keep, flip

OpenStudy (anonymous):

2.666666666666/1.5=1.7777777777777333

OpenStudy (anonymous):

2 2/3 = 8/3 1 1/2 = 3/2 (8/3) / (3/2) = (8/3) * (2/3) = 16/9 So the answer is 16/9 or 1.777...

OpenStudy (anonymous):

Some people complicate things too much. ^

OpenStudy (anonymous):

Thanks guys :)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

@OrangeMaster It was "complicated" because I showed how to do it without a calculator or long division.

OpenStudy (anonymous):

Almost everybody on here can use a calculator, because most computers have one already installed. It was easier to do with a calculator.

OpenStudy (anonymous):

@OrangeMaster I'm not critizing your approach, just so you know. I just feel that it's important to know how to work with such simple fractions without a calculator.